OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -9) and a directrix y = 9. explain how to do this please?

OpenStudy (anonymous):

@deepconcept did your face fall asleep on the keyboard lol

OpenStudy (anonymous):

let a arbitrary point P (x,y) on parabola. from definition of parabola, distance of point P (x,y) from focus S (0, -9) is equal to its distance from directrix.(say MN) then, SP= perpendicular distance of P from MN.... proceed now.....

OpenStudy (anonymous):

i'm confused @deepconcept

OpenStudy (anonymous):

can you walk me through this?

OpenStudy (anonymous):

(y-9)^2 = (x-0)^2 + (y+9)^2

OpenStudy (anonymous):

solving , we get x^2 = -36y

OpenStudy (anonymous):

and to solve for y we get y=-x^2/36?

OpenStudy (anonymous):

it's not the standard form of parabola.

OpenStudy (anonymous):

well that's my answer choice..

OpenStudy (anonymous):

all my answer choices are solved for either y or y^2..

OpenStudy (anonymous):

have you got the answer?